40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10422     Entry time: Fri Aug 22 03:55:45 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Xarm PDH fine, Yarm PDH/ALS needs work 

[Rana, Jenne, EricQ]

We did several things tonight.  First, a list (so I can remember them all), and then some details.

(1) Jiggled ETMY SUS cables, removed kicks.

(2) Locked X and Y ALS, looked at POX, POY as out of loop sensors.

(3) Measured stuff (?) at the Yend.

(4) Reconnected REFL DC to SR560.

(5) Attempted CARM offset reduction.


Item 1:

When Rana and I started locking this evening, we saw (as Q has been witnessing for a while now) the ETMY kick a lot.  However, it seemed to be kicking even more than usual.  Since Q had been down at the end station recabling things, we wondered if a SUS-related cable got bumped.  Rana went down to the end and pushed all the cables into their receptacles.  One of the last sets that he pushed was the satellite box.  We didn't have walkie-talkie communication, but the DC offset of the ETMY oplevs changed just a minute or two before he returned to the control room.  So, we guess that it was the satellite box cables that were loose.  Unfortunately, there is no clear way to strain relieve them, which is why they can so often be troublesome.  Anyhow, the ETMY hasn't kicked since.

Item 2:

We locked the arms with ALS.  We saw that the POX signal was about 20% of the full pk-pk height of the PDH signal, so it's mostly within the linear range, but not entirely.  It is what it is, however, and we took measurements assuming that it's okay.  I calibrated POX by putting an excitation onto ETMX, and matching the height of the peak in POX and BEATX_FINE_PHASE_OUT_HZ.

Q and Rana had also [remembered / put in / something] a digital readback for the end green PDH error point.  Q went down to the end and gave me a number of 2600 Hz/V for the err mon port of the PDH board, which is what is connected to the ADC.  With that and 20/2^16 V/cts, I had a calibration of 0.8 Hz/ct. 

What we see in this plot is that the green end PDH is not the limiting noise for the POX out of loop measurement of the residual arm motion.  Also, in the multi-color metrology paper, Fig 7 (which is posted in the control room), we see at about a little over 1 Hz a ratio of about 4.5 between the residual motion and the AUX PDH error signal.  In today's plot, I see a ratio of about 20.  I infer from this that the green PDH for the Xarm is fine, and that we may want to re-look at the ALS digital loop, but we should leave the X PDH alone.

Here is the Xarm plot:

Xend_ErrorPointMeasurements.pdf

Q took the data for the Yarm plot, so hopefully he can give it to us in the morning.  What we did notice was that the noise was much worse for the Yarm.  This prompted Item 3, measuring the loop.

Item 3:

Q and Rana went down to the Yend and measured some things.  They came back, and said that they hadn't changed anything in analog while they were down there.  One thing that Q did note was that we have almost 90 degrees of phase margin (since it's a 1/f loop), and about 10 dB of gain margin, above the UGF.  So, we're in good shape for being able to try triggering the boost on the PDH box.  Q will give us more notes on this work, as well as plots, in the morning.

Item 4:

At some point, I remembered that Q and Gabriele had repurposed the SR560 that we had been using for the REFLDC input to the common mode board.  So, Q went and put it back, so that REFL DC goes into the SR560, and so does a DAC channel so that we can remotely set the offset.  The A-B output goes to the REFL11I whitening channel, since real REFL11I goes into the input of the CM board.  I think that today, the SR 560 was left at a gain of 1.

Item 5:

We decided to carry on and try to reduce the CARM offset some.  An annoyance is that the Yarm still has pretty significant low-frequency noise, but the idea is that if we can get over to the sqrtInvTrans signals, it will be fine.

So, we didn't get much farther than we had in the past, but it was nice to get there at all again.  I ran the carm_cm_up script (many times).  One of the times, all I wanted to do was see how much I could reduce the CARM offset.  CARM was on sqrtInvTrans, DARM was on ALS diff, and I was able to get the arm powers up to about 2.5.  I don't know why I lost lock.  The sqrtInv signals should be good until at least arm powers of 20 or so. 

I was able to see the REFL DC dip, but only a teensy tiny bit.  It went down by maybe 1 count.  Q suggested looking at how deep it could get while leaving CARM and DARM both on ALS, and setting both offsets to 0.  We were seeing arm flashes of about 50 counts, and REFL DC went from 0 to -800.  So, I wasn't seeing much of a REFL dip, but it was definitely there when I went to arm powers of 2ish.

We tried looking at different sqrtInv options for DARM, and haven't come to any real conclusion.  In the plot below, we are looking at a swept sine between DARM_IN1 (ALSdiff) and either MC_IN1 0.3*(sqrtInvX - sqrtInvY) or SRCL_IN1 (TRX - TRY / sqrt(TRX + TRY) ):

DARM_ALSdiff_vs_sqrtInv.pdf

 


We have a few things to add to the to-do list:

* Put UGF servos for LSC loops in place.

* Implement UGF "servos" (per Koji's suggested method) for phase trackers.

* Write a lockloss script that is run by the ALS watch scripts - print a PDF of error and control signals for every lockloss, and save it somewhere.

* Fix up Ygreen modematching on the PSL table.  The X green spot is quite similar on the camera to the corresponding PSL green spot.  However the Y green spot is not at all the same as its PSL green spot. 

 

ELOG V3.1.3-